OpenStudy (anonymous):

restriction of (x^2+3x+2)/(x^2-2x-3) * (x^2+4x+3)/(x+2)

OpenStudy (anonymous):

x ≠ -3 x ≠ 3 x ≠ 1 x ≠ -2 I'm getting different results so I'm lost.. I'm getting a result of x plus or minus -1, -2 and 3

hero (hero):

\(\left(\dfrac{x^2 + 3x + 2}{x^2 - 2x - 3}\right) \left(\dfrac{x^2 + 4x + 3}{x + 2}\right)\)

OpenStudy (anonymous):

Yes....

hero (hero):

Now...without doing any work at all, what can x not equal?

hero (hero):

We know that \(x + 2 \ne 0\) right?

OpenStudy (anonymous):

mhm

hero (hero):

So what can x not equal?

OpenStudy (anonymous):

I understand the extra help but I just need to know which of my results are correct... I'm getting two that match my answer choices but I can only click one

OpenStudy (anonymous):

It's not giving a direct answer imo since I've done the work already

OpenStudy (anonymous):

that's what I have marked, thanks.

OpenStudy (anonymous):

Nah, I said only I can pick one

hero (hero):

Well, factors of one can cancel...so maybe we should try to factor this one.

OpenStudy (anonymous):

Huh...? I said I can only pick one answer which is -2 o.o

hero (hero):

\(\left(\dfrac{(x+2)(x+1)}{(x-3)(x+1)}\right) \left(\dfrac{(x+3)(x+1)}{x + 2}\right)\) Yes, I know, however, if factors can cancel, then we'll be able to reduce it to just one answer choice.

hero (hero):

\(\left(\dfrac{\cancel{(x+2)}\cancel{(x+1)}}{(x-3)\cancel{(x+1)}}\right) \left(\dfrac{(x+3)(x+1)}{\cancel{x + 2}}\right)\)

hero (hero):

Now we're left with just: \(\dfrac{(x+3)(x+1)}{x-3}\)

hero (hero):

So \(x - 3 \ne 0\) and \(x \ne 3\)

OpenStudy (anonymous):

Well... How did I get -1 then? o_o does that mean it's changed to 3 and not -2?

hero (hero):

Since factors of one cancel, we can reduce the fraction to a simplified form and find what x definitely cannot equal.

OpenStudy (anonymous):

so.. x cannot equal -2 and 3? it looks that way due to how you wrote it (I have another question like this)

hero (hero):

Sorry, but if we can cancel factors, then they are no longer restrictions. \(x \ne 3\) is the correct choice.

OpenStudy (anonymous):

I suppose that makes sense :O the other one is (5x^2-10x-15)/(x^2-9) divided by (6x+12)/(x+2) I've gotten multiple results as well. choices are: x ≠ -2 x ≠ -3 x ≠ 3 x ≠ 9

OpenStudy (anonymous):

is it B, -3?

hero (hero):

You basically do the same thing as before which is reduce it to simplest form first, then figure out the restriction.

hero (hero):

@whalexnuker show your work

OpenStudy (anonymous):

I see you typing- I got it, thanks ;o

hero (hero):

What do you mean, you got it? What work did you do to get it?

OpenStudy (anonymous):

I got the answer.. I submitted the test and got full credit o.o that's why the question was closed earlier. I just haven't strayed from the question, lol